# Unifews: Unified Entry-Wise Sparsification for Efficient Graph Neural Network

### v1

@@ Abstract.
Graph Neural Networks (GNNs) have shown promising performance in various graph learning tasks, but at the cost of resource-intensive computations. The primary overhead of GNN update stems from graph propagation and weight transformation, both involving operations on graph-scale matrices. Previous studies attempt to reduce the computational budget by leveraging graph-level or network-level sparsification techniques, resulting in downsized graph or weights. In this work, we propose Unifews, which unifies the two operations in an entry-wise manner considering individual matrix elements, and conducts joint edge-weight sparsification to enhance learning efficiency. The entry-wise design of Unifews enables adaptive compression across GNN layers with progressively increased sparsity, and is applicable to a variety of architectural designs with on-the-fly operation simplification. Theoretically, we establish a novel framework to characterize sparsified GNN learning in view of a graph optimization process, and prove that Unifews effectively approximates the learning objective with bounded error and reduced computational load. We conduct extensive experiments to evaluate the performance of our method in diverse settings. Unifews is advantageous in jointly removing more than $90\%$ of edges and weight entries with comparable or better accuracy than baseline models. The sparsification offers remarkable efficiency improvements including 1020×10-20$\times$10 - 20 × matrix operation reduction and up to 100×100$\times$100 × acceleration in graph propagation time for the largest graph at the billion-edge scale.
1. Introduction
@@
Graph Neural Networks (GNNs) have undergone extensive development for learning from graph-structure data and have achieved remarkable performance on a variety of tasks (Kipf and Welling, 2017; Hamilton et al., 2017; Veličković et al., 2018; Zhang and Chen, 2018). The power of GNNs is mainly attributed to their message-passing scheme that each node in the graph aggregates information from its neighbors, namely graph propagation, and then updates its own representation accordingly by trainable weights as feature transformation. Iteratively performing the two-stage scheme enables the model to learn from the graph topology inclusively.
@@
Despite their success, canonical GNNs are recognized for their high resource consumption, especially when scaling up to large graphs (Wu et al., 2021). The performance bottleneck arises from the sequence of matrix multiplications integral to message passing stages, i.e., graph propagation and feature transformation, which alternatively apply to the node representation and lead to computational complexities directly proportional to the numbers of edges and nodes in the graph, respectively (Chen et al., 2020). Hence, the essence of improving GNN learning efficiency lies in reducing the number of operations associated with graph diffusion and model weights (Liu et al., 2022; Zhang et al., 2023a).
Empirically, a diverse range of strategies have been explored for saving computational cost by sparsifying the graph as well as the model. In efforts to simplify graph computation, prior research utilizes graph sparsification techniques (Liu et al., 2019). This typically entails removing graph components based on either predetermined (Zheng et al., 2020; Zeng et al., 2021; Liu et al., 2023b) or learned (Li et al., 2021, 2022; Jin et al., 2022) criteria. However, the persistence of a static graph through all propagation hops causes a dilemma: the topology may become overly coarse, thereby omitting crucial information for GNN feature extraction, or the graph structure may be under-sparsified and model efficiency is scarcely improved.
@@
Another direction is to exploit compression on the model architecture by integrating neural network pruning approaches (Deng et al., 2020a), which gradually sparsifies weight elements during GNN training based on performance assessments (Zhou et al., 2021; Chen et al., 2021; You et al., 2022; Wang et al., 2023). Regardless of the improved flexibility, their scalability for large graphs remains limited, as full-graph computation is still necessary at certain stages of the pipeline. Moreover, there is a noticeable gap in the theoretical interpretation of GNN sparsification despite its practical utility. Existing works are either constrained to the layer-agnostic graph approximation (Chen et al., 2018; Zou et al., 2019), or only provide straightforward representation error analysis (Srinivasa et al., 2020; Liu et al., 2023b). The impact of simplified graphs on the GNN learning process, particularly through multiple layers, remains inadequately addressed.
To mitigate the above drawbacks of current compression solutions, we propose to rethink both graph propagation and feature transformation from an entry-wise perspective, i.e., examining individual matrix elements. As illustrated in Figure 1, the two operations are performed by employing the diffusion and weight matrices, respectively. Directly sparsifying entries in these two matrices enables us to reduce the number of entry operations. The entry-wise mechanism functions adaptively, that it not only simplifies the current computation dynamically, but also transmits the outcome to succeeding GNN layers and facilitates further sparsification.
@@
Figure 1. (a) Within one hop of aggregation to the center node, messages from neighbors with larger diffusion values are associated with greater importance. (b) Unifews jointly applies sparsification to both graph propagation and feature transformation stages in each GNN layer. Color intensity of an entry denotes its relative magnitude. Unimportant entries of the diffusion and weight matrices are pruned in order to reduce computational operations.
In this paper, we propose Unifews, a Unified entry-wise sparsi-fication framework for GNN graph and weights. Theoretically, we establish a high-level framework for characterizing the unified sparsification. To our knowledge, we are the first to associate joint sparsification with the GNN optimization process. Our analysis elucidates the effect of sparsification across iterative GNN layers, which differs from the previous approach focusing on a particularized metric for approximation. Based on the theory, we derive a quantitative guarantee for Unifews on the error bound with respect to the graph learning objective, and demonstrate that the amount of reduced operations is at least linear to the sparsity.
In practice, Unifews is capable of jointly removing unnecessary operations in both graph propagation and feature transformation towards improved model efficiency. The entry-wise strategy is adeptly integrated into matrix computations on the fly without incurring additional overhead. The progressive sparsification across layers further fosters an increase in sparsity for both stages, which reciprocally augments their efficiency. Consequently, Unifews is powerful in conducting personalized propagation as well as transformation for each node and each layer, effectively balancing the operational reduction while retaining precision. The theoretical and implementation framework of Unifews is applicable to a broad family of message-passing GNNs, covering the representative models of both iterative and decoupled architectures.
Our main contributions are outlined as the following:
  • We propose Unifews as a unified GNN sparsification, which employs a dedicated entry-wise scheme on both graph connections and model weights. Unifews is adaptable to common GNN designs to reduce graph-scale operations and alleviate computational cost.
  • We develop novel theoretical analysis and guarantee of Unifews approximation through the comprehensive lens of GNN learning process. Unifews enhances sparsity by adaptively simplifying the current operation and accumulating it in future layer updates.
  • We experimentally evaluate Unifews with a wide range of GNN architectures and graph datasets. Our method outperforms state-of-the-art GNN compressions by reaching over $90\%$ of joint sparsity without compromising accuracy. On the billion-scale graph papers100M, Unifews is able to boost the computation time by 100×100$\times$100 ×.
2. Related Work
@@
GNNs with Graph Simplification. The motivation of graph simplification is to modify the graph components, i.e., nodes and edges, and augment the learning process (Liu et al., 2022). Among them, a large portion of works utilize graph sparsification for accuracy improvement, typically through deleting edges. NeuralSparse (Zheng et al., 2020) and LSP (Kosman et al., 2022) drop edges based on $k$-neighbor metric for each node and layer-wise local graph topology, respectively, while AdaptiveGCN (Li et al., 2021) and SGCN (Li et al., 2022) choose to implement sparsity by learnable edge features and graph adjacency. In comparison, FastGAT (Srinivasa et al., 2020) and DSpar (Liu et al., 2023b) relate graph modification to spectral sparsification mainly as an attempt to boost efficiency. FastGAT eliminates connections in the GAT (Veličković et al., 2018) model by calculating the Effective Resistance (Spielman and Srivastava, 2011) but at the expense of high computational time. DSpar employs a loose approximation and successfully applies to larger datasets. Graph coarsening describes another simplification approach that reduces the graph size by contracting nodes into subsets. GraphZoom (Deng et al., 2020b), GOREN (Cai et al., 2021), and Huang et al. (2021) explore various coarsening schemes and enhance training scalability, while GCond (Jin et al., 2022) alternatively adopts an analogous condensation strategy. Most of the simplification techniques, however, conduct graph modification in a one-time and layer-agnostic manner, resulting in limited flexibility and utility under the trade-off between efficiency and efficacy. Additionally, we distinguish these methods from graph sampling (Chen et al., 2018; Huang et al., 2018; Zou et al., 2019; Zeng et al., 2020; Rong et al., 2020; Fang et al., 2023b) that does not produce a deterministic sparsified graph and is usually constrained for solely enhancing the training process.

### v2

# Unifews: You Need Fewer Operations for Efficient Graph Neural Networks

@@ Abstract
Graph Neural Networks (GNNs) have shown promising performance, but at the cost of resource-intensive operations on graph-scale matrices. To reduce computational overhead, previous studies attempt to sparsify the graph or network parameters, but with limited flexibility and precision boundaries. In this work, we propose Unifews, a joint sparsification technique to unify graph and weight matrix operations and enhance GNN learning efficiency. The Unifews design enables adaptive compression across GNN layers with progressively increased sparsity, and is applicable to a variety of architectures with on-the-fly simplification. Theoretically, we establish a novel framework to characterize sparsified GNN learning in view of the graph optimization process, showing that Unifews effectively approximates the learning objective with bounded error and reduced computational overhead. Extensive experiments demonstrate that Unifews achieves efficiency improvements with comparable or better accuracy, including 1020×10-20$\times$10 - 20 × matrix operation reduction and up to 100×100$\times$100 × acceleration for graphs up to billion-edge scale. Our code is available at: https://github.com/gdmnl/Unifews.
Graph Neural Networks, Acceleration, Graph Spectral Theory, Graph Sparsification
1 Introduction
@@
Graph Neural Networks (GNNs) have undergone extensive development for learning from graph-structure data and have achieved remarkable performance (Kipf & Welling, 2017; Hamilton et al., 2017; Veličković et al., 2018). The power of GNNs is mainly attributed to the message-passing scheme containing two alternative operations, namely graph propagation and feature transformation, which aggregates neighborhood information in the graph and updates representation by trainable weights.
@@
Despite their success, canonical GNNs are recognized for their high resource consumption, especially when scaling up to large graphs (Wu et al., 2021). The performance bottleneck arises from the connection with graph size, as both graph propagation and feature transformation can be viewed as multiplications on graph-scale matrices, and lead to computational complexities proportional to the numbers of edges and nodes in the graph, respectively (Chen et al., 2020). Hence, the essence of improving GNN learning efficiency lies in reducing the number of operations associated with graph diffusion and model weights (Liu et al., 2022; Zhang et al., 2023a).
@@
In efforts to reduce the graph part of computation, prior GNN graph sparsification techniques (Liu et al., 2019) typically remove graph components based on either predetermined (Zheng et al., 2020; Zeng et al., 2021; Liu et al., 2023b) or learned (Li et al., 2021, 2022; Jin et al., 2022) criteria. Another direction is to exploit compression on the model architecture by integrating network pruning approaches (Deng et al., 2020a), which gradually sparsifies weight elements during GNN training (Zhou et al., 2021; Chen et al., 2021; You et al., 2022; Wang et al., 2023b). Both approaches, however, suffer from limited flexibility due to the static graph structure throughout all propagation iterations: the topology may become overly coarse, thereby omitting crucial information for GNN feature extraction, or may be under-sparsified and model efficiency is scarcely improved. Moreover, there is a noticeable gap in the theoretical interpretation of GNN sparsification. Existing works are either constrained to the layer-agnostic graph approximation (Chen et al., 2018; Zou et al., 2019), or only provide straightforward representation error analysis (Srinivasa et al., 2020; Liu et al., 2023b). The impact of simplified graphs on the GNN learning process, particularly through multiple layers, remains inadequately addressed.
To mitigate the drawbacks of current GNN compression, we propose Unifews, a Unified entry-wise sparsi-fication framework for GNN graph and weights. As illustrated in Figure 1, by considering the two stages in GNN as matrix operations, graph and model sparsification can be unified in an entry-wise manner, i.e., manipulating individual elements during matrix multiplication. We then establish a theoretical framework to quantify the impact of Unifews on precision and efficiency. By bridging the graph learning process in the spectral domain, our analysis elucidates the effect of joint sparsification across iterative GNN layers in a holistic manner, which differs from previous approaches focusing on a particularized metric for approximation.
In practice, Unifews is capable of reducing operations in the two-stage GNN message-passing and improving efficiency. The simple but effective strategy is powerful in conducting flexible computation for each node and each GNN layer, effectively balancing the operational reduction while retaining precision. The theoretical and implementation framework of Unifews is applicable to a broad family of GNNs, covering the representative models of both iterative and decoupled architectures.
In summary, (1) we propose Unifews for unifying GNN graph and model sparsification from an entry-wise perspective to reduce graph-scale operations and alleviate computational cost; (2) we develop novel theoretical framework bridging graph spectral filtering and sparsification as approximation; (3) we experimentally demonstrate that Unifews outperforms existing GNN compressions by reaching over $90\%$ of joint sparsity without compromising accuracy. On the billion-scale graph papers100M, Unifews is able to boost the computation time by 100×100$\times$100 ×.
2 Preliminaries and Related Works
2.1 GNN Architecture
Graph Notation.
Consider a graph $\mathcal{G}=\langle\mathcal{V},\mathcal{E}\rangle$ with $n=|\mathcal{V}|$ nodes and $m=|\mathcal{E}|$ edges. The neighborhood set of a node $u\in\mathcal{V}$ is $\mathcal{N}(u)=\{v\mid(u,v)\in\mathcal{E}\}$, and its degree $d(u)=|\mathcal{N}(u)|$. The self-looped adjacency matrix is $\bar{\bm{A}}\in\mathbb{R}^{n\times n}$ and the diagonal degree matrix is $\bm{D}=\operatorname{diag}(d(u_{1}),d(u_{2}),\cdots,d(u_{n}))$. The normalized adjacency and Laplacian matrices are $\tilde{\bm{A}}=\bm{D}^{-1/2}\bar{\bm{A}}\bm{D}^{-1/2}$ and $\tilde{\bm{L}}=\bm{I}-\tilde{\bm{A}}$, respectively.
Iterative GNN.
Iterative models such as GCN (Kipf & Welling, 2017) take the node attribute $\bm{X}\in\mathbb{R}^{n\times f}$ as input, and recurrently update by applying the graph diffusion matrix $\bm{T}$. The representation matrix $\bm{H}_{(l+1)}$ of the $(l+1)$-th layer is updated as $\bm{H}_{(l+1)}=\sigma\left(\bm{T}\bm{H}_{(l)}\bm{W}_{(l)}\right)$, where $l\in\{0,\cdots,L-1\}$, and $\sigma(\cdot)$ denotes the activation function. It implies two consecutive steps, that graph propagation computes the embedding $\bm{P}_{(l)}=\bm{T}\bm{H}_{(l)}$, and feature transformation multiplies the learnable weight $\bm{W}_{(l)}$. For example, the diffusion matrix of GCN is $\bm{T}=\tilde{\bm{A}}$.
Decoupled GNN.
This variant simplifies GNN architecture by separating the propagation from iterative updates (Wu et al., 2019; Gasteiger et al., 2019; Wang et al., 2021a). The graph-related propagation is computed in advance as the embedding matrix $\bm{P}_{(l+1)}=\bm{T}_{(l)}\cdot\bm{P}_{(l)}$ for $l\in\{0,\cdots,L-1\}$, and the transformation is as simple as an $L$-layer MLP that $\bm{H}_{(l+1)}=\sigma\left(\bm{H}_{(l)}\bm{W}_{(l)}\right)$, where the boundary conditions are $\bm{P}_{(0)}=\bm{X}$ and $\bm{H}_{(0)}=\bm{P}_{(L)}$. As an exemplar, the diffusion matrices for SGC (Wu et al., 2019) and APPNP (Klicpera et al., 2019) are $\bm{T}=\tilde{\bm{A}}$ and $\bm{T}=\alpha(1-\alpha)\tilde{\bm{A}}$ according to their respective designs.
@@ 2.1 GNN Architecture
Figure 1: (a) Graph propagation messages from neighbors with larger diffusion values are associated with greater importance. (b) Unifews jointly applies sparsification to both GNN stages. Unimportant entries of the diffusion and weight matrices are pruned to reduce operations.
(a) Propagation Personalization
(b) Joint Compression
(c) Unifews (on Iterative GNN)
Figure 2: Comparison of GNN pipelines between conventional simplification techniques and our Unifews framework. (a) Personalized propagation iteratively simplifies the node-dependent graph diffusion but is only applicable to decoupled GNNs. (b) Joint model compression can sparsify both graph and weight, whereas the same diffusion is utilized across all layers. (c) Unifews implements entry-wise sparsification of both graph and weights for each layer with increased sparsity.
2.2 Taxonomy of GNN Sparsification
Iterative Graph Sparsification.
Graph sparsification methods aim to augment the iterative message-passing and improve accuracy (Liu et al., 2022). NeuralSparse (Zheng et al., 2020) and LSP (Kosman et al., 2022) drop edges based on topological properties, while AdaptiveGCN (Li et al., 2021) and SGCN (Li et al., 2022) implement learnable graph connections. FastGAT (Srinivasa et al., 2020) and DSpar (Liu et al., 2023b) specifically examine spectral patterns. However, these techniques modify the graph in a one-time and layer-agnostic manner, resulting in limited flexibility for balancing efficiency and efficacy.
Decoupled Propagation Personalization.
A particularized branch emerges for simplifying decoupled GNNs (Zhang et al., 2021; Huang et al., 2023; Liao et al., 2023b; Zeng et al., 2021). Since the graph operation is isolated, it is more flexible for fine-grained modification, which replaces the graph-level message-passing with separate maneuver on the edge connection for each propagation hop as in Figure 2(a). Our Unifews is akin to this idea, but with a better suitability for both iterative and decoupled propagations.
Architectural and Joint Compression.
A series of research take the network architecture into account. Their general pipeline, as outlined in Figure 2(b), involves a model compressor and a static graph sparsifier. Some works treat the two stages independently (Zhou et al., 2021; Liu et al., 2023a), while others exploit joint updates (You et al., 2022; Sui et al., 2024; Chen et al., 2021; Wang et al., 2023b, a). One critical drawback of these methods lies in the additional bottleneck for learning the compression strategy, which still incurs full-scale computation and are less efficient. A thorough discussion regarding other efficient GNNs such as graph sampling can be found in Appendix A.
3 Graph Smoothing under Sparsification
In this section, we first relate GNN learning with the graph smoothing process, which enables us to characterize the process by our novel approximation bound. Generally, a broad scope of both iterative and decoupled GNNs can be interpreted by a spectral graph smoothing process (Zhu et al., 2021). We adopt the optimization framework as:
Definition 3.1 (Graph Laplacian Smoothing (Ma et al., 2021)).
$\bm{p}^{\ast}=\operatorname*{arg\,min}_{\bm{p}}\mathcal{L},\quad\mathcal{L}=\|\bm{p}-\bm{x}\|^{2}+c\cdot\bm{p}^{\top}\bm{L}\bm{p},$ (1)
@@ 3 Graph Smoothing under Sparsification
In Equation 1, $\bm{L}$ is the general Laplacian matrix, as normalization only causes a difference in coefficient. The first term of $\mathcal{L}$ reflects the closeness to the input signal, i.e., node attributes representing their identity. The second term is associated with the graph structure, acting as a regularization that constrains the representation values of connected node pairs to be similar. The closed-form solution $\bm{p}^{\ast}=(\bm{I}+c\bm{L})^{-1}\bm{x}$ can be inferred when the derivative $\partial\mathcal{L}/\partial\bm{p}=0$. However, note that it is prohibitive to directly acquire the converged solution due to the inverse calculation on large graph-scale matrix. Hence, GNN models employ an iterative approach to learn the representation under this framework with varying architectural designs. For example, GCN convolution corresponds to $c=1$.
Next, we introduce graph sparsifiers as an approximation process of graph smoothing. In order to measure the extend of approximation and its impact on learning outcomes, we consider the spectral similarity:
Definition 3.2 ($\epsilon$-Spectral Similarity).
The approximate Laplacian matrix $\hat{\bm{L}}$ is said to be $\epsilon$-spectrally similar to the raw Laplacian matrix $\bm{L}$ if:
$\bm{x}^{\top}(\hat{\bm{L}}-\epsilon\bm{I})\bm{x}\leq\bm{x}^{\top}\bm{L}\bm{x}\leq\bm{x}^{\top}(\hat{\bm{L}}+\epsilon\bm{I})\bm{x},\;\forall\bm{x}\in\mathbb{R}^{n}.$ (2)
Graph approximation satisfying Definition 3.2 can be regarded as spectral sparsification (Spielman & Srivastava, 2011; Batson et al., 2013), which identifies operations that maintains certain spectral properties such as eigenvalues during modification. Compared to the common multiplicative spectral similarity (Sadhanala et al., 2016; Calandriello et al., 2018; Charalambides & Hero, 2023), Definition 3.2 possesses an additive tolerance, which allows for manipulating specific entries of $\bm{L}$ and suits our scenario.
Then, we are able to characterize the graph smoothing problem for sparsified graphs. Under edge modifications, we intend to bound the optimization goal under approximation:
Lemma 3.3 (Approximate Graph Laplacian Smoothing).
Given two graphs $\mathcal{G}=\langle\mathcal{V},\mathcal{E}\rangle$ and $\hat{\mathcal{G}}=\langle\mathcal{V},\hat{\mathcal{E}}\rangle$, where $\hat{\mathcal{E}}$ is the sparsified edge set. When the Laplacian $\hat{\bm{L}}$ of $\hat{\mathcal{G}}$ is $\epsilon$-similar to $\bm{L}$ of $\mathcal{G}$, the solution $\hat{\bm{p}}^{\ast}$ to the problem Equation 1 w.r.t $\hat{\bm{L}}$ is called an $\epsilon$-approximation of the solution $\bm{p}^{\ast}$ w.r.t. $\bm{L}$ , and:
$\|\hat{\bm{p}}^{\ast}-\bm{p}^{\ast}\|\leq c\epsilon\|\bm{p}^{\ast}\|.$ (3)
The proof of Lemma 3.3 can be found in Section B.1. It establishes a novel interpretation for characterizing the smoothing procedure under a sparsified graph, that if a sparsifier complies with the spectral similarity Definition 3.2, it is capable of effectively approximating the iterative graph optimization and achieving a close output with bounded error. Compared to approximation bounds specifying feature values in previous GNN sparsification theories (Srinivasa et al., 2020; Zhang et al., 2023b), our analysis highlights the impact of graph sparsifier throughout the holistic learning process, which enjoys better expressiveness and suitability.
4 Unifews Algorithm and Theory
This section presents our Unifews framework by respectively developing its application to decoupled and iterative GNNs, where graph and weight sparsification are separately performed in the former architecture, and are combined in the latter. Our key theoretical insights are summarized as:
  • Bridging sparsification and smoothing: Unifews can be described as an spectral sparsifier. Its sparsification strength is determined by the threshold of entry removal.
  • Multi-layer bounds: Unifews provides an approximation to the graph smoothing optimization. Its representation error to the original objective is effectively bounded.
  • Improvements on efficiency and efficacy: Unifews is advantageous in enhancing efficiency, mitigating the over-smoothing issue, and facilitating enhanced joint sparsity.
4.1 Unifews as Spectral Sparsification
Intuition: Entry Values Denote Importance in Graph Computation.
As depicted in Figure 2(b), conventional graph sparsification methods for GNNs only provide fixed and uniform graph-level adjustments through the entire learning process. This lack of flexibility hinders their performance when employing to the recurrent update design with multiple GNN layers. Contrarily, node-wise models in Figure 2(a) aim to personalize the graph-based propagation, but come with additional explicit calculations and impaired runtime efficiency. We are hence motivated to design a sparsification approach that (1) enables modifications in a fine-grained manner to further simplify the computation; and (2) can be seamlessly integrated into the matrix operation with negligible overhead.
To this end, we first focus on the isolated graph propagation stage and express the propagation of decoupled models in an entry-wise manner on node $u$ as:
$\small\bm{p}_{(l+1)}[u]=\sum_{v\in\mathcal{N}(u)}\tau_{(l)}[u,v]=\sum_{v\in\mathcal{N}(u)}\bm{T}_{(l)}[u,v]\cdot\bm{p}_{(l)}[v],$ (4)
where the propagation message $\tau_{(l)}[u,v]$ from node $v$ to $u$ is regarded as an entry. As illustrated in Figure 1(a), in a single hop of GNN diffusion, edges carrying propagation messages exert varying impacts on neighboring nodes, whose importance is dependent on the graph topology. Messages with minor significance are usually considered redundant and can be eliminated to facilitate sparsity. From the perspective of matrix computation, this is achieved by omitting the particular message $\tau[u,v]$ in current aggregation based on an importance indicator such as its magnitude.
Edge Pruning for Single Layer.
In order to perform pruning on entry $\tau[u,v]$, the diffusion matrix $\bm{T}$ can be utilized to record the pruning information, which is exactly the concept of graph sparsification. Given a universal magnitude threshold $\delta_{a}$, zeroing out entries $|\tau[u,v]|<\delta_{a}$ is equivalent to sparsifying $\bm{T}$ with a node-wise threshold $\delta^{\prime}_{a}$ as:
$\hat{\bm{T}}[u,v]=\operatorname{thr}_{\delta^{\prime}_{a}}\big{(}\bm{T}[u,v]\big{)}\cdot\bm{T}[u,v],\quad\delta^{\prime}_{a}=\delta_{a}/|\bm{p}[v]|,$ (5)
where the pruning function with an arbitrary threshold $\delta$ is $\operatorname{thr}_{\delta}(x)=1$ if $|x|>\delta$, and $\operatorname{thr}_{\delta}(x)=0$ otherwise. Sparsification by Equation 5 has two concurrent effects: for graph computation, messages with small magnitudes are prevented from propagating to neighbors and composing the output representation; for graph topology, corresponding edges are removed from the diffusion process.
Next, we show that edge pruning in Equation 5 for one layer can be considered as a spectral sparsification following Definition 3.2, which enables us to derive its approximation bound. Practically, the target sparsity $\eta_{a}\in[0,1]$ is usually determined by the realistic application. Let $\hat{\mathcal{E}}$ be the edge set achieved by Unifews sparsification and the corresponding Laplacian is $\hat{\bm{L}}$, there is $\eta_{a}=1-|\hat{\mathcal{E}}|/m$. We first derive the following lemma as a general guarantee associating sparsification and spectral similarity:
Lemma 4.1.
Given graph $\mathcal{G}$ and embedding $\bm{p}$, let $\hat{\mathcal{E}}$ be the edge set achieved by graph sparsification with threshold $\delta_{a}$. The sparsified Laplacian matrix $\hat{\bm{L}}$ is $\epsilon$-similar to $\bm{L}$ when $q_{a}\delta_{a}\leq\epsilon\|\bm{p}\|$, where $q_{a}$ is the number of edges removed.
The proof of Lemma 4.1 is detailed in Section B.2. To derive more specific bounds of $q_{a}$, we need to examine the actual distribution of entry values. Regarding the assumption on edge distribution, we particularly investigate the scale-free graph, which is common in realistic large-scale graphs (Clauset et al., 2009). For input attribute values, we consider the Gaussian distribution, which is also commonly used for depicting the feature distribution of neural networks as an extension of the central limit theorem (Deshpande et al., 2018; Bojchevski & Günnemann, 2018). With these two assumptions, the following theorem relates the threshold and approximation bound with sparsification:
Theorem 4.2 (Bound for Unifews).
Given a graph $\mathcal{G}$, embedding $\bm{p}$, and required edge sparsity $\eta_{a}$, the threshold $\delta_{a}$ can be decided by $\delta_{a}=C(1-\eta_{a})^{-t}$, and the sparsified Laplacian $\hat{\bm{L}}$ is $\epsilon$-similar to $\bm{L}$ with the approximation bound:
$\epsilon=O\left(\eta_{a}(1-\eta_{a})^{-t}\right),$ (6)
where $C$ and $t$ are positive constants. $C>0$ is determined by the embedding values $\|\bm{p}\|$, and $10<t<$ 1 depicts the degree distribution.
The proof of Theorem 4.2 and specific derivations of $C$ and $t$ can be found in Section B.3. The significance of Theorem 4.2 lies in bridging the sparsification technique and Laplacian smoothing with a specific precision bound for the first time to the best of our knowledge. Its implication is that, the adaptive sparsification by Unifews for a single layer qualifies as a graph spectral sparsifier bounded by $\epsilon$. Its pivotal parameter, i.e., the sparsification threshold, and its approximation bound can be determined once given the sparsity $\eta_{a}$. A sparsity closer to $1$ results in a larger threshold value as well as a loose error guarantee.
4.2 Unifews for Graph Propagation
Thanks to the adaptive property of the entry-wise pruning scheme, we are able to further perform fine-grained and gradual sparsification across propagation layers. Intuitively, a message deemed of minor significance in the current propagation is unlikely to propagate further and influence more distant nodes in subsequent layers. Unifews hence offers progressively increasing sparsity throughout GNN layers.
As depicted in Figure 2(c), Unifews iteratively applies sparsification to each layer, and the pruned diffusion matrix $\hat{\bm{T}}_{(l)}$ is inherited to the next layer for further edge removal. Denote the edge set corresponding to $\hat{\bm{T}}_{(l)}$ as ^$\hat{\mathcal{E}}_{(l)}$, there is ^$\hat{\mathcal{E}}_{(l)}\subseteq\hat{\mathcal{E}}_{(l-1)}\subseteq\cdots\subseteq\hat{\mathcal{E}}_{(0)}\subseteq\mathcal{E}$, which indicates a diminishing number of operations for deeper layers. Consequently, Equation 4 is modified as the following to represent entry-level diffusion on the sparsified graph:
$\small\hat{\bm{p}}_{(l+1)}[u]=\sum_{v\in\mathcal{N}(u)}\hat{\bm{T}}_{(l)}[u,v]\cdot\bm{p}_{(l)}[v]=\sum_{v\in\hat{\mathcal{N}}_{(l)}(u)}\tau_{(l)}[u,v],$ (7)
where the neighborhood composed by remaining connections is $\hat{\mathcal{N}}_{(l)}(u)=\{v\mid(u,v)\in\hat{\mathcal{E}}_{(l)}\}$.